OpenStudy (anonymous):

find the equation with the following zeros: -1,4,-5i

OpenStudy (anonymous):

I'm guess you are listing three separate zeros. Where (x+1)=0 is one zero. Where (x-4)=0 is another zero. And (x+5i)=0 is a third zero. Note that complex zeros come in pairs, so there would be a fourth zero such as (x-5i)=0 Is this what you meant?

OpenStudy (anonymous):

Woops didn't see the comma there...marshallinwashington has nearly the right factors but still, given 0 restrictions on coefficients, (x-5i) is not necessarily a root. There are infinitely many such equations however.